Can someone help me solve for y:
5(4x + 3) <= -1/2y + 5
20x + 15 <= -1/2y + 5
20x + 15 - 5 <= -1/2y
20x + 10 <= -1/2y
I get lost after this.

Answer by rapaljer(4671)

What you have done so far looks right.
20x + 10 <= -1/2y

Now multiply both sides by -2, which of course reverses the inequality sign:
-2(20x + 10) >= -2(-1/2y)
-40x - 20 >= y

You probably don't want y on the right side so switch sides, but when you do this reverses the inequality again:
-40x - 20 >= y
y<=-40x-20

Is that what you wanted to do??

R^2 at SCC